OpenStudy (anonymous):

Which is the equation of the line that contains points (0, 5) and (5, 8)?

geerky42 (geerky42):

Do you know Point Slope Form? \(y - y_1 = m(x -x_1)\) where \((x_1,y_1)\) is point. you would need to use slope formula \(m = \dfrac{y_2 - y_1}{x_2 - x_1}\) to find slope